Nineteen Measures of Solvency and Profitability

The comparative financial statements of Blige Inc. are as follows. The market price of Blige Inc. common stock was $64 on December 31, 2016.

Blige Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 2016 and 2015
    2016     2015
Retained earnings, January 1 $4,221,375 $3,569,725
Add net income for year 976,800 731,200
Total $5,198,175 $4,300,925
Deduct dividends
On preferred stock $12,600 $12,600
On common stock 66,950 66,950
Total $79,550 $79,550
Retained earnings, December 31 $5,118,625 $4,221,375


Blige Inc.
Comparative Income Statement
For the Years Ended December 31, 2016 and 2015
    2016     2015
Sales $6,910,980 $6,358,100
Sales returns and allowances 34,380 22,350
Sales $6,876,600 $6,335,750
Cost of goods sold 2,286,360 2,103,450
Gross profit $4,590,240 $4,232,300
Selling expenses $1,667,160 $2,020,260
Administrative expenses 1,420,180 1,186,500
Total operating expenses 3,087,340 3,206,760
Income from operations $1,502,900 $1,025,540
Other income 79,100 65,460
$1,582,000 $1,091,000
Other expense (interest) 472,000 260,000
Income before income tax $1,110,000 $831,000
Income tax expense 133,200 99,800
Net income $976,800 $731,200


Blige Inc.
Comparative Balance Sheet
December 31, 2016 and 2015
    Dec. 31, 2016     Dec. 31, 2015
Assets
Current assets
Cash $1,358,870 $927,060
Temporary investments 2,056,660 1,536,280
Accounts receivable (net) 1,182,600 1,109,600
Inventories 890,600 686,200
Prepaid expenses 257,083 185,410
Total current assets $5,745,813 $4,444,550
Long-term investments 1,976,300 23,815
Property, plant, and equipment (net) 7,080,000 6,372,000
Total assets $14,802,113 $10,840,365
Liabilities
Current liabilities $1,853,488 $1,438,990
Long-term liabilities
Mortgage note payable, 8%, due 2021 $2,650,000 $0
Bonds payable, 8%, due 2017 3,250,000 3,250,000
Total long-term liabilities $5,900,000 $3,250,000
Total liabilities $7,753,488 $4,688,990
Stockholders' Equity
Preferred $0.7 stock, $50 par $900,000 $900,000
Common stock, $10 par 1,030,000 1,030,000
Retained earnings 5,118,625 4,221,375
Total stockholders' equity $7,048,625 $6,151,375
Total liabilities and stockholders' equity $14,802,113 $10,840,365

Required:

Determine the following measures for 2016, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

ROUND ALL TO 2 DECIMALS

1. Number days' sales in inventory (Round to 2 decimals) ______

2. Ratio of fixed assets to long-term liabilities (Round to 2 decimals) ______

3. Ratio of liabilities to stockholders' equity (Round to 2 decimals) ______

4. Number of times interest charges are earned (Round to 2 decimals) ______

5. Number of times preferred dividends are earned (Round to 2 decimals) ______

6. Ratio of sales to assets (Round to 2 decimals) ______

7. Rate earned on total assets (Round to 2 decimals) ______

8. Rate earned on stockholders' equity (Round to 2 decimals) ______

9. Rate earned on common stockholders' equity (Round to 2 decimals) ______

10. Earnings per share on common stock (Round to 2 decimals) ______

11. Price-earnings ratio (Round to 2 decimals) ______

12. Dividends per share of common stock (Round to 2 decimals) ______

13. Dividend yield (Round to 2 decimals) ______
Answer #1

1. Number days' sales in inventory = (Inventory/Cost of goods sold) * 365 days

=(890600/2286360) * 365 days

= 142 Days Approx.

2. Ratio of fixed assets to long-term liabilities  = Fixed Assets/Long Term Liabilities

= (1976300+7080000)/5900000

=1.53 times

3. Ratio of liabilities to stockholders' equity = Total Liabilities/stockholders' equity

=7753488/7048625

=1.10 times

4. Number of times interest charges are earned = Earnings Before Interest & Tax/Interest Expense

=1582000/472000

=3.35 times

5. Number of times preferred dividends are earned = Dividend on preferred stock/preferred stock

=12600/900000

=0.014 times

6. Ratio of sales to assets = Net Sales (Total Sales-Sales Return)/Total Assets

=6876600(6910980-34380)/14802113

=0.46 times

7. Rate earned on total assets = Net Income/Average Total Assets(Opening+closing Total Assets)/2

=976800/(10840365+14802113)/2

=976800/12821239

=0.0762 times or 7.62%

8. Rate earned on stockholders' equity = Net Income/Average Stockholders' Equity(opening+closing stockholders' Equity)/2

=976800/(7048625+6151375)/2

=976800/6600000

=0.148 or 14.8%

9. Rate earned on common stockholders' equity=

(Net Income - Dividend on Preferred stock)/Average Common Stock(opening + closing common stockholders' equity)

=(976800-12600)/(1030000+1030000)/2

=964200/1030000

=0.94 times or 94%

10. Earnings per share on common stock = (Net Income - Dividend on Preferred stock)/Common Stock(in numbers)

=(976800-12600)/103000

=9.36 times

11. Price-earnings ratio = Market Price per share/ Earnings per share

=64(given)/9.36

=6.84 times

12. Dividends per share of common stock = Dividend to common stock/common stock (in number)

=66950/103000

=0.65 times

13. Dividend yield = Dividends per share of common stock/Market Price per share

=(0.65/64) * 100

= 1.02 %